Firm Beta has invested $20 billion in building its plant. To finance this investment, Beta borrowed $14 billion and raised the remaining funds by issuing new shares of stock. Suppose there are currently 100 million shares of stock outstanding. Investors are very excited about Beta’s growth, and thus think that the stock price of $75 is reasonable. Suppose these are Beta’s only assets.

(1) What’s the book value of shareholders’ equity?

(2) Assume the market value of debt outstanding is also $14 billion. What’s the market value of total assets?

(3) Comparing the book value and market value of assets, which value is larger? Is it surprising?
Answer #1
1] Book value of shareholders' equity = 20-14 = $               6.00 billion
2] Market value of total assets = MV of debt+MV of equity = 14+0.1*75 = $            21.50 billion
3] Market value is larger. No, it is not surprising as, the
MV of equity would be higher for a profit making
company.